OpenStudy (anonymous):

to find the area of a triangle do you add or multiply all 3 sides?

OpenStudy (akashdeepdeb):

NONE!! :D There are specific formulae for different triangles. :D Area of a right triangle is = height times base over 2 Area of equilateral = sqrt3 * side * side over 4 Area of Isosceles = Altitude times base/2 Look up some videos on youtube for more!! :D Understood? :)

OpenStudy (akashdeepdeb):

I am not getting your question that well @Bbambino1976 . :/

OpenStudy (akashdeepdeb):

To find PERMETER YOU ADD ALL SIDES!! :D

OpenStudy (anonymous):

Forgive me, it"s an equilateral. So

OpenStudy (akashdeepdeb):

aND What are you finding? Area or permeter? :D

OpenStudy (anonymous):

I'm trying to find the area, with 1600 on a side, with vertices at 3 points?

OpenStudy (akashdeepdeb):

http://www.mathwords.com/a/area_equilateral_triangle.htm

OpenStudy (anonymous):

133.333 how does that look?

OpenStudy (akashdeepdeb):

Are you sure the side is 1600? Because 133.333333 seems to be incorrect. :|

OpenStudy (anonymous):

the question says its 1600 km on a side,hey how did you know that?

OpenStudy (anonymous):

it says the Bermuda triangle is a equilateral triangle 1600k on a side with vertices in Bermuda, Puerto Rico, and Florida coast. Assuming it is flat, what is the approximate area?

OpenStudy (anonymous):

a2 = (a/2)2 + h2 a2 = a2/4 + h2 a2 − a2/4 = h2 4a2/4 − a2/4 = h2 3a2/4 = h2 h = √(3a2/4) h = (√(3)×a)/2 Area = (base × h)/2 base × h = (a × √(3)×a)/2 = (a2× √(3))/2 Dividing by 2 is the same as multiplying the denominator by 2. Therefore, the formula is equilateral-triangle-formula-image

OpenStudy (anonymous):

so do i take 1600 square it then divide it by two?

OpenStudy (anonymous):

no that can't be right it's only 20